# Orientation double covers of non-orientable Lefschetz fibrations

### v1

@@ Abstract
    for an auto-diffeomorphism $\omega:N_{g}\to N_{g}$.
@@
In this subsection, we recall some properties of orientation double covering. More details can be found in [13] for example.
@@
    For each $q\in N$,the inverse image $\Theta^{-1}(q)$ is the finite set consisting of two points;
@@
Let $M$ be an $m$-dimensional smooth manifold with a smooth atlas $\{(U_{\lambda},\varphi_{\lambda})\}_{\lambda\in\Lambda}$ and $\mathrm{ob}(T_{p}M)$ the set of all ordered bases for the tangent space at $p\in M$. We say that two ordered bases are equivalent if the change-of-basis matrix has positive determinant. This gives an equivalence relation $\sim$’ on $\mathrm{ob}(T_{p}M)$, and hence, the quotient set $\mathrm{ob}(T_{p}M)/\sim$ of $\mathrm{ob}(T_{p}M)$ is defined. $\mathrm{ob}(T_{p}M)/\sim$ consists of exactly two elements, which are called local orientations of $M$ around $p$. Now, let $\widetilde{M}$ be a set
@@
where $x_{1}^{\lambda},\cdots,x_{m}^{\lambda}$ is local coordinates on $U_{\lambda}$, and $\Big{[}\Big{(}\frac{\partial}{\partial x_{1}^{\lambda}}\Big{|}_{p},\cdots,\frac{\partial}{\partial x_{m}^{\lambda}}\Big{|}_{p}\Big{)}\Big{]}\in\mathrm{ob}(T_{p}M)/\sim$ is the equivalence class of the ordered basis $\Big{(}\frac{\partial}{\partial x_{1}^{\lambda}}\Big{|}_{p},\cdots,\frac{\partial}{\partial x_{m}^{\lambda}}\Big{|}_{p}\Big{)}$ for $T_{p}M$. Then we see that $\pi$ is a double covering map, especially an orientation double covering map and $\{({U_{\lambda}}^{+},{\varphi_{\lambda}}^{+}),({U_{\lambda}}^{-},{\varphi_{\lambda}}^{-})\}_{\lambda\in\Lambda}$ provides a smooth atlas of $\widetilde{M}$. From now on, to simplify notation, let $o_{p}$ (resp. $o_{p}$) denote $\Big{[}\Big{(}\frac{\partial}{\partial x_{1}^{\lambda}}\Big{|}_{p},\cdots,\frac{\partial}{\partial x_{m}^{\lambda}}\Big{|}_{p}\Big{)}\Big{]}$ (resp. $\Big{[}\Big{(}\frac{\partial}{\partial x_{1}^{\lambda}}\Big{|}_{p},\cdots,\frac{\partial}{\partial x_{m}^{\lambda}}\Big{|}_{p}\Big{)}\Big{]}$). We remark that each of orientations of $T_{(p,+o_{p})}\widetilde{M}$ and $T_{(p,-o_{p})}\widetilde{M}$ are exactly specified by
@@
In particular, for various discussions, we need the next lemma.
@@
From now on, for a smooth manifold $Q$, $\mathrm{Int}(Q)$ and $\partial Q$ denote the interior and the boundary of $Q$, respectively. In addition, for a smooth map $F$ between two manifolds, we write the set of all critical points of $F$ as $\mathrm{Crit}(F)$ and the image $F(\mathrm{Crit}(F)$ as $\Delta_{F}$.
@@
We call $X$ and $\Sigma$ the total space and the based space of a Lefschetz fibration, respectively.
@@
Let $f:X\to\Sigma$ be a Lefschetz fibration. Since Definition 2.6 (LF3) and compactness of X, it follows that $\mathrm{Crit}(f)$ is the finite set. Moreover, we can assume that $f|_{\mathrm{Crit}(f)}$ is the injection, i.e., each singular fiber of $f$ has only one critical point by perturbing $f$ slightly.
@@
We first show (LF1) and (LF2). Since $\pi$ is the local diffeomorphism, we observe, by the excision theorem, that $H_{4}(\widetilde{X},\widetilde{X}-\{(x,o_{x})\})\cong H_{4}(X,X-\{x\})$ for any $x\in X$. Thus we have
@@
Also, $\pi$ is the submersion since $\pi$ is obviously the smooth covering map, and hence, we see that
$\begin{split}\mathrm{Crit}(f\circ\pi)&=\{(x,o_{x})\in\widetilde{X}\thickspace|\thickspace x\in\mathrm{Crit}(f)\}.\\ &=\pi^{-1}(\mathrm{Crit}(f)).\end{split}$
@@
Next, we show (LF3). Fix $x\in\mathrm{Crit}(f)$. Take charts $(U,\varphi)$ and $(V,\phi)$ about $x$ and $f(x)$, respectively, such that they satisfy the condition (LF3). From the discussion of Example 2.3, $(U^{\dagger},\varphi^{\dagger})$ is a chart about $(x,{\dagger}o_{x})\in\widetilde{X}$ for each $\dagger$$\in\{$+,-\}, and we then get
@@
for all $(z,w)\in\varphi^{\dagger}(U^{\dagger})=\varphi(U)\subset\mathbb{C}^{2}$. Therefore, combining this with the fact that $\varphi^{+}$ and $\varphi^{-}$ is orientation preserving and orientation reversing, respectively, it follows that $f\circ\pi$ is an achiral Lefschetz fibration which has the same number of positive critical points and negative critical points.
@@
In Proposition 1.1, another proof about the genus of $f\circ\pi:\widetilde{X}\to\Sigma:=\Sigma_{g^{\prime}}^{b}$ is given by using an Euler characteristic on the condition that we assume that $2g^{\prime}+b\neq 2$. Suppose that $f$ has $n$ singular fibers. Then,
@@
because $X$ is the smooth manifold, i.e., the CW complex and the degree of the covering map $\pi:\widetilde{X}\to X$ is two. Moreover, according to the proof of Proposition 1.1 and Remark 2.7, the number of singular fibers of $f\circ\pi$ is $2n$. So we have
@@
is diffeomorphic to the circle with center $0\in\mathbb{C}$ and radius $\sqrt{|\varepsilon|}$ if $\varepsilon\neq 0$ and the set consisting of the single point $0\in\mathbb{C}^{2}$ if $\varepsilon=0$. Let $f:X\to\Sigma$ be a Lefschetz fibration. Fix $y_{0}\in\Delta_{f}$ and $y_{1}\in\Sigma-\Delta_{f}$, and let $x_{0}\in\mathrm{Crit}(f)\cap f^{-1}(y_{0})$. We can take a smooth embedding $\ell:[0,1]\to\Sigma$ such that $y_{0}=\ell(0)$ and $y_{t}:=\ell(t)\in\Sigma-\Delta_{f}$ $(t\in(0,1])$. In addition, choose complex coordinate charts $(U,\varphi)$ and $(V,\psi)$ about $x_{0}$ and $y_{0}$, respectively, such that they satisfy the condition (LF3). Now, for a sufficiently small fixed $\delta\in(0,1]$, we define a simple closed curve as
@@
for any $x\in f^{-1}(y_{t})$, where $T(f^{-1}(\ell([\delta,1])))$ represents the tangent bundle of $f^{-1}(\ell([\delta,1]))$. For each $x\in f^{-1}(y_{\delta})$, we then have an integral curve $c_{x}$ of $W$ starting at $x$, and therefore, we get the diffeomorphism $\Upsilon:f^{-1}(y_{\delta})\to f^{-1}(y_{1})$ given by $\Upsilon(x):=c_{x}(1-\delta)$ for all $x\in f^{-1}(y_{\delta})$. Furthermore, we obtain, by using the diffeomorphism $\Upsilon$, the simple closed curve $\Upsilon(C)\subset f^{-1}(y_{1})$. Note that the isotopy class of $\Upsilon(C)$ is uniquely determined by the homotopy class of $\ell$ with fixed endpoints. In the context of discussion above, each of $\Upsilon(C)$ and its isotopy class is said to be the vanishing cycle of $f$ determined by $\ell$.
@@
Let $f:X\to\Sigma$ be a genus-$g$ non-orientable Lefschetz fibration, and we will use the same symbols which are used in the above discussion (for instance $x_{0},\ell,C,\Upsilon$, etc) for $f$. It follows from Proposition 1.1 and its proof that $f\circ\pi:\widetilde{X}\to\Sigma$ is an achiral Lefschetz fibration of genus-(g1)(g-1) and a chart about $(x_{0},{\dagger}o_{x_{0}})\in\mathrm{Crit}(f\circ\pi)\cap(f\circ\pi)^{-1}(y_{0})$ with the condition (LF3) is given by $(U^{{\dagger}},\varphi^{{\dagger}})$ for each $\dagger$$\in\{$+,-\}. As above, for each $\dagger$$\in\{$+,-\}, we can construct the simple closed curve
@@
Also, we can define a diffeomorphism $\widetilde{\Upsilon}:(f\circ\pi)^{-1}(y_{\delta})\to(f\circ\pi)^{-1}(y_{1})$ using an integral curve $\widetilde{c}_{(x,o_{x})}$ of $\widetilde{W}$ starting at $(x,o_{x})\in(f\circ\pi)^{-1}(y_{\delta})$ to be $\widetilde{\Upsilon}((x,o_{x})):=\widetilde{c}_{(x,o_{x})}(1-\delta)$ for all $(x,o_{x})\in(f\circ\pi)^{-1}(y_{\delta})$. Note that
@@
In the process of proving Proposition 3.1, we also find vanishing cycles of a non-orientable Lefschetz fibration $f$ are two-sided. Indeed, since $\pi|_{U^{{\dagger}}}:U^{{\dagger}}\to U$ is the diffeomorphism for $\dagger$$\in\{$+,-\}, it follows that $C=\pi|_{U^{{\dagger}}}(C^{{\dagger}})$ is two-sided, and thus, $\Upsilon(C)$ is also two-sided.
@@
Let $\mathcal{M}(S_{g})$ be the set of all isotopy classes of auto-diffeomorphisms of $S_{g}$, here if $S_{g}=\Sigma_{g}$, we add a condition that auto-diffeomorphisms are orientation preserving. This forms a group and said to be the mapping class group of $S_{g}$. Note here that we give the operation on $\mathcal{M}(S_{g})$ by $\omega]\cdot[\omega^{\prime}]:=[\omega^{\prime}\circ\omega]$ for $\omega],[\omega^{\prime}]\in\mathcal{M}(S_{g})$. It is well-known that $\mathcal{M}(\Sigma_{g})$ is generated by Dehn twists (cf. [2]). Moreover, $t_{c_{1}}t_{c_{2}}=t_{c_{2}}t_{c_{1}}$ if $c_{1}$ and $c_{2}$ are disjoint simple closed curves in $\Sigma_{g}$ and $t_{\omega(c)}=\Omega^{-1}t_{c}\Omega$ for $\Omega=[\omega]\in\mathcal{M}(\Sigma_{g})$ are also well-known. On the other hand, it is well-known fact that we need not only Dehn twists but also isotopy classes of other diffeomorphisms which are called crosscap slides to generate $\mathcal{M}(N_{g})$ (cf. [11] and [12]). The following relations hold in $\mathcal{M}(N_{g})$:
@@
and
$t_{\omega(c);\omega_{*}(\theta_{c})}=\Omega^{-1}t_{c;\theta_{c}}\Omega$ (4.2)
for $\Omega=[\omega]\in\mathcal{M}(N_{g})$, where $\theta_{c}$ is the opposite orientation with respect to $\theta_{c}$ and $\omega_{*}(\theta_{c})$ means the orientation of $\nu(\omega(c))$ induced by $\theta_{c}$.
@@
For $g\geqq 3$, the map $\eta:\mathcal{M}(N_{g})\to\mathcal{M}(\Sigma_{g-1})$ defined by $\eta([\omega]):=[\widetilde{\omega}$] is the injective homomorphism, where $\widetilde{\omega}$ is exactly one orientation preserving lift of $\omega$.
@@
Let $f:X\to\Sigma$ be a genus-$g$ Lefschetz fibration. Fix $y\in\Sigma-\Delta_{f}$ and a diffeomorphism $\Xi:f^{-1}(y)\to S_{g}$, here $\Xi$ is required to preserve orientation if $S_{g}=\Sigma_{g}$. Taking a smooth loop $a:[0,1]\to\Sigma-\Delta_{f}$, the trivial bundle $a^{*}f|_{f^{-1}(\Sigma-\Delta_{f})}:a^{*}f^{-1}(\Sigma-\Delta_{f})\to[0,1]$ is induced from the fiber bundle $f|_{f^{-1}(\Sigma-\Delta_{f})}:f^{-1}(\Sigma-\Delta_{f})\to\Sigma-\Delta_{f}$ with the fiber $S_{g}$ by $a$. Let $\Phi:[0,1]\times S_{g}\to a^{*}f^{-1}(\Sigma-\Delta_{f})$ be a trivialization map of $a^{*}f|_{f^{-1}(\Sigma-\Delta_{f})}$ with $\Phi(0,p)=(0,\Xi^{-1}(p))$ for all $p\in S_{g}$, and for each $t\in[0,1]$, define a map $\Phi_{t}:S_{g}\to f^{-1}(y)$ as $\Phi_{t}(p):=proj_{2}\circ\Phi(t,p)\thickspace(p\in S_{g})$, where $proj_{2}:[0,1]\times f^{-1}(\Sigma-\Delta_{f})\to f^{-1}(\Sigma-\Delta_{f})$ is the second projection. Then we get a well-defined homomorphism $\rho_{f}:\pi_{1}(\Sigma-\Delta_{f},y)\to\mathcal{M}(S_{g})$ which is defined by $\rho_{f}([a]):=[\Xi\circ\Phi_{1}]$ for all $a]\in\pi_{1}(\Sigma-\Delta_{f},y)$. This is called a monodromy representation of $f$.
@@
We say that each of $a_{i}$ and its homotopy class is the meridian loop with respect to $\mu_{i}$. It is well-known that $\rho_{f}([a_{i}])$ is equal to the Dehn twist about the vanishing cycle of $f$ determined by $\mu_{i}$. Here, we note a few things about this. First, consider the case of $f$ being orientable. In this case, the Dehn twist about the vanishing cycle determined by $\mu_{i}$ is right handed (resp. left handed) if the critical point of $f$ belonging to $\mathrm{Crit}(f)\cap f^{-1}(y_{i})$ is positive (resp. negative). Next, consider the case of $f$ being non-orientable. From Remark 3.2, notice that the Dehn twist about each of vanishing cycles of $f$ can be defined. In this case, whether the Dehn twist about the vanishing cycle determined by $\mu_{i}$ is right handed or left handed depends on an orientation of a tubular neighborhood of the vanishing cycle because of the relation (4.1).
@@
It is easily seen that $\Sigma-\Delta_{f\circ\pi}=\Sigma-\Delta_{f}$ holds. By applying Lemma 2.5, for the diffeomorphism $\Xi:f^{-1}(y)\to N_{g}$, we get a diffeomorphism $\widetilde{\Xi}:(f\circ\pi)^{-1}(y)\to\Sigma_{g-1}$ such that the diagram (2.1) commutes. Take the pull-back $a^{*}(f\circ\pi)|_{(f\circ\pi)^{-1}(\Sigma-\Delta_{f})}:a^{*}(f\circ\pi)^{-1}(\Sigma-\Delta_{f})\to[0,1]$ of the fiber bundle $(f\circ\pi)|_{(f\circ\pi)^{-1}(\Sigma-\Delta_{f})}:(f\circ\pi)^{-1}(\Sigma-\Delta_{f})\to\Sigma-\Delta_{f}$ by $a$. Note that $(id_{[0,1]}\times\pi)|_{a^{*}(f\circ\pi)^{-1}(\Sigma-\Delta_{f})}:a^{*}(f\circ\pi)^{-1}(\Sigma-\Delta_{f})\to a^{*}f^{-1}(\Sigma-\Delta_{f})$ is an orientation double covering map since $a^{*}f^{-1}(\Sigma-\Delta_{f})$ is a submanifold of $0,1]\times f^{-1}(\Sigma-\Delta_{f})$ and $a^{*}(f\circ\pi)^{-1}(\Sigma-\Delta_{f})=\pi^{-1}(a^{*}f^{-1}(\Sigma-\Delta_{f}))$. Lemma 2.5 hence gives a unique orientation preserving diffeomorphism $\widetilde{\Phi}:[0,1]\times\Sigma_{g-1}\to a^{*}(f\circ\pi)^{-1}(\Sigma-\Delta_{f})$ for the diffeomorphism $\Phi:[0,1]\times N_{g}\to a^{*}f^{-1}(\Sigma-\Delta_{f})$ so that the upper rectangle of the diagram

### v2

@@ Abstract
    for an auto-diffeomorphism $\omega:N_{g}\to N_{g}$ of the non-orientable, closed, connected, genus-$g$ smooth surface $N_{g}$.
Here, for a two-sided simple closed curve $c$ on $N_{g}$, $t_{c;\theta_{c}}$} denotes the right handed Dehn twist about $c$ considering an orientation $\theta_{c}$ of a tubular neighborhood of $c$. Further details are provided in Subsection 4.1.
@@
In this subsection, we recall some properties of orientation double covering maps. More details can be found in [13] for example.
@@
    For each $q\in N$,the inverse image $\Theta^{-1}(q)$ is a finite set consisting of two points;
@@
Let $M$ be an $m$-dimensional smooth manifold with a smooth atlas $\{(U_{\lambda},\varphi_{\lambda})\}_{\lambda\in\Lambda}$ and $\mathrm{ob}(T_{p}M)$ the set of all ordered bases for the tangent space $T_{p}M$ at $p\in M$. We say that two ordered bases are equivalent if the change-of-basis matrix has positive determinant. This gives an equivalence relation $\sim$’ on $\mathrm{ob}(T_{p}M)$, and hence, the quotient set $\mathrm{ob}(T_{p}M)/\sim$ of $\mathrm{ob}(T_{p}M)$ is defined. The set $\mathrm{ob}(T_{p}M)/\sim$ consists of exactly two elements, which are called local orientations of $M$ around $p$. Now, let $\widetilde{M}$ be the set
@@
$U_{\lambda}^{-}:=\Big\{(p,o_{p})\in\widetilde{M}\thickspace|\thickspace p\in U_{\lambda},o_{p}=-\Big[\Big(\frac{\partial}{\partial x_{1}^{\lambda}}\Big|_{p},\cdots,\frac{\partial}{\partial x_{m}^{\lambda}}\Big|_{p}\Big)\Big]\Big\},$
@@
where $x_{1}^{\lambda},\cdots,x_{m}^{\lambda}$ is the local coordinate on $U_{\lambda}$, and $\Big[\Big(\frac{\partial}{\partial x_{1}^{\lambda}}\Big|_{p},\cdots,\frac{\partial}{\partial x_{m}^{\lambda}}\Big|_{p}\Big)\Big]\in\mathrm{ob}(T_{p}M)/\sim$ is the equivalence class of the ordered basis $\Big(\frac{\partial}{\partial x_{1}^{\lambda}}\Big|_{p},\cdots,\frac{\partial}{\partial x_{m}^{\lambda}}\Big|_{p}\Big)$ for $T_{p}M$. Then we see that $\pi$ is a double covering map, especially an orientation double covering map and $\{({U_{\lambda}}^{+},{\varphi_{\lambda}}^{+}),({U_{\lambda}}^{-},{\varphi_{\lambda}}^{-})\}_{\lambda\in\Lambda}$ provides a smooth atlas of $\widetilde{M}$. From now on, to simplify notation, let $o_{p}$ (resp. $o_{p}$) denote $\Big[\Big(\frac{\partial}{\partial x_{1}^{\lambda}}\Big|_{p},\cdots,\frac{\partial}{\partial x_{m}^{\lambda}}\Big|_{p}\Big)\Big]$ (resp. $\Big[\Big(\frac{\partial}{\partial x_{1}^{\lambda}}\Big|_{p},\cdots,\frac{\partial}{\partial x_{m}^{\lambda}}\Big|_{p}\Big)\Big]$). We remark that orientations of $T_{(p,+o_{p})}\widetilde{M}$ and $T_{(p,-o_{p})}\widetilde{M}$ are exactly specified by
@@
In particular, for various discussions in this paper, we need the next lemma.
@@
From now on, for a smooth manifold $Q$, $\mathrm{Int}(Q)$ and $\partial Q$ denote the interior and the boundary of $Q$, respectively. In addition, for a smooth map $F$ between two manifolds, we write the set of all critical points of $F$ as $\mathrm{Crit}(F)$ and the image $F(\mathrm{Crit}(F))$ as $\Delta_{F}$.
@@
We call $X$ and $\Sigma$ the total space and the based space of a Lefschetz fibration $f:X\to\Sigma$, respectively.
@@
Let $f:X\to\Sigma$ be a Lefschetz fibration. Since the condition (LF3) and compactness of X, it follows that $\mathrm{Crit}(f)$ is a finite set. Moreover, we can assume that $f|_{\mathrm{Crit}(f)}$ is an injection, i.e., each singular fiber of $f$ has only one critical point by perturbing $f$ slightly.
@@
We first show that $f\circ\pi$ satisfies the conditions (LF1) and (LF2). Since $\pi$ is a local diffeomorphism, we observe, by the excision theorem, that $H_{4}(\widetilde{X},\widetilde{X}-\{(x,o_{x})\})\cong H_{4}(X,X-\{x\})$ for any $x\in X$. Thus we have
@@
Since $\pi$ is a smooth covering map, it is in particular a submersion. Hence, we see that
$\begin{split}\mathrm{Crit}(f\circ\pi)&=\{(x,o_{x})\in\widetilde{X}\thickspace|\thickspace x\in\mathrm{Crit}(f)\}\\ &=\pi^{-1}(\mathrm{Crit}(f)),\end{split}$
@@
Next, we show that $f\circ\pi$ satisfies the condition (LF3). Fix $x\in\mathrm{Crit}(f)$. Take charts $(U,\varphi)$ and $(V,\phi)$ about $x$ and $f(x)$, respectively, such that they satisfy the condition (LF3). From the discussion of Example 2.3, for $(U,\varphi)$, $(U^{\dagger},\varphi^{\dagger})$ is a chart about $(x,{\dagger}o_{x})\in\widetilde{X}$ for each $\dagger$$\in\{$+,-\}, and we then get
@@
for all $(z,w)\in\varphi^{\dagger}(U^{\dagger})=\varphi(U)\subset\mathbb{C}^{2}$. Therefore, combining this with the fact that $\varphi^{+}$ and $\varphi^{-}$ are orientation preserving and orientation reversing, respectively, it follows that $f\circ\pi$ is an achiral Lefschetz fibration which has the same number of positive critical points and negative critical points.
@@
In Proposition 1.1, an alternative proof of the genus of $f\circ\pi:\widetilde{X}\to\Sigma:=\Sigma_{g^{\prime}}^{b}$ is given using the Euler characteristic of a Lefschetz fibration, under the assumption that $2g^{\prime}+b\neq 2$. Suppose that $f$ has $n$ singular fibers. Then
@@
because $X$ is a smooth manifold, i.e., a CW complex and the degree of the covering map $\pi:\widetilde{X}\to X$ is two. Moreover, according to the proof of Proposition 1.1 and Remark 2.7, the number of singular fibers of $f\circ\pi$ is $2n$. So we have
@@
is diffeomorphic to the circle with center $0\in\mathbb{C}$ and radius $\sqrt{|\varepsilon|}$ if $\varepsilon\neq 0$ and equals to the set consisting of the single point $0\in\mathbb{C}^{2}$ if $\varepsilon=0$. Let $f:X\to\Sigma$ be a Lefschetz fibration. Fix $y_{0}\in\Delta_{f}$ and $y_{1}\in\Sigma-\Delta_{f}$, and let $x_{0}\in\mathrm{Crit}(f)\cap f^{-1}(y_{0})$. We can take a smooth embedding $\ell:[0,1]\to\Sigma$ such that $y_{0}=\ell(0)$ and $y_{t}:=\ell(t)\in\Sigma-\Delta_{f}$ $(t\in(0,1])$. In addition, choose complex coordinate charts $(U,\varphi)$ and $(V,\psi)$ about $x_{0}$ and $y_{0}$, respectively, such that they satisfy the condition (LF3). Now, for a sufficiently small fixed $\delta\in(0,1]$, we define a simple closed curve as
@@
for any $x\in f^{-1}(y_{t})$, where $T(f^{-1}(\ell([\delta,1])))$ represents the tangent bundle of $f^{-1}(\ell([\delta,1]))$. For each $x\in f^{-1}(y_{\delta})$, we then have an integral curve $c_{x}$ of $W$ starting at $x$, and therefore, we get the diffeomorphism $\Upsilon:f^{-1}(y_{\delta})\to f^{-1}(y_{1})$ given by $\Upsilon(x):=c_{x}(1-\delta)$ for all $x\in f^{-1}(y_{\delta})$. Furthermore, we obtain, by using the diffeomorphism $\Upsilon$, the simple closed curve $\Upsilon(C)\subset f^{-1}(y_{1})$. Note that the isotopy class of $\Upsilon(C)$ is uniquely determined by the homotopy class of $\ell$ with fixed endpoints. In the context of the discussion above, the isotopy class of $\Upsilon(C)$ is said to be the vanishing cycle of $f$ determined by $\ell$.
@@
Let $f:X\to\Sigma$ be a genus-$g$ non-orientable Lefschetz fibration, and for this, we will use the same symbols as in the above discussion (for instance, $x_{0},\ell,C,\Upsilon$, etc). It follows from Proposition 1.1 and its proof that $f\circ\pi:\widetilde{X}\to\Sigma$ is an achiral Lefschetz fibration of genus-(g1)(g-1) and a chart about $(x_{0},{\dagger}o_{x_{0}})\in\mathrm{Crit}(f\circ\pi)\cap(f\circ\pi)^{-1}(y_{0})$ with the condition (LF3) is given by $(U^{{\dagger}},\varphi^{{\dagger}})$ for each $\dagger$$\in\{$+,-\}. As above, for each $\dagger$$\in\{$+,-\}, we can construct the simple closed curve
@@
We can define a diffeomorphism $\widetilde{\Upsilon}:(f\circ\pi)^{-1}(y_{\delta})\to(f\circ\pi)^{-1}(y_{1})$ using an integral curve $\widetilde{c}_{(x,o_{x})}$ of $\widetilde{W}$ starting at $(x,o_{x})\in(f\circ\pi)^{-1}(y_{\delta})$ to be $\widetilde{\Upsilon}((x,o_{x})):=\widetilde{c}_{(x,o_{x})}(1-\delta)$ for all $(x,o_{x})\in(f\circ\pi)^{-1}(y_{\delta})$. Note that
@@
In the process of proving Proposition 3.1, we also find vanishing cycles of a non-orientable Lefschetz fibration are two-sided. Indeed, since $\pi|_{U^{{\dagger}}}:U^{{\dagger}}\to U$ is a diffeomorphism for $\dagger$$\in\{$+,-\}, it follows that $C=\pi|_{U^{{\dagger}}}(C^{{\dagger}})$ is two-sided, and thus, $\Upsilon(C)$ is also two-sided.
@@
Let $\mathcal{M}(S_{g})$ be the set of isotopy classes of auto-diffeomorphisms of $S_{g}$, here if $S_{g}=\Sigma_{g}$, we add a condition that auto-diffeomorphisms are orientation preserving. This forms a group and said to be the mapping class group of $S_{g}$. Note here that we give the operation on $\mathcal{M}(S_{g})$ by $\omega]\cdot[\omega^{\prime}]:=[\omega^{\prime}\circ\omega]$ for $\omega],[\omega^{\prime}]\in\mathcal{M}(S_{g})$. It is well-known that $\mathcal{M}(\Sigma_{g})$ is generated by Dehn twists (cf. [2]). Moreover, $t_{c_{1}}t_{c_{2}}=t_{c_{2}}t_{c_{1}}$ if $c_{1}$ and $c_{2}$ are disjoint simple closed curves in $\Sigma_{g}$ and $t_{\omega(c)}=\Omega^{-1}t_{c}\Omega$ for $\Omega=[\omega]\in\mathcal{M}(\Sigma_{g})$ are also well-known. On the other hand, it is well-known fact that we need not only Dehn twists but also isotopy classes of other diffeomorphisms which are called crosscap slides to generate $\mathcal{M}(N_{g})$ (cf. [11] and [12]). The following relations hold in $\mathcal{M}(N_{g})$:
@@
and for $\Omega=[\omega]\in\mathcal{M}(N_{g})$,
$t_{\omega(c);\omega_{*}(\theta_{c})}=\Omega^{-1}t_{c;\theta_{c}}\Omega,$ (4.2)
where $\theta_{c}$ is the opposite orientation with respect to $\theta_{c}$ and $\omega_{*}(\theta_{c})$ means the orientation of $\nu(\omega(c))$ induced by $\theta_{c}$.
@@
For $g\geqq 3$, the map $\eta:\mathcal{M}(N_{g})\to\mathcal{M}(\Sigma_{g-1})$ defined by $\eta([\omega]):=[\widetilde{\omega}$] is an injective homomorphism, where $\widetilde{\omega}$ is exactly one orientation preserving lift of $\omega$.
@@
Let $f:X\to\Sigma$ be a genus-$g$ Lefschetz fibration. Fix $y\in\Sigma-\Delta_{f}$ and a diffeomorphism $\Xi:f^{-1}(y)\to S_{g}$, here $\Xi$ is assumed to preserve orientation if $S_{g}=\Sigma_{g}$. For the fiber bundle $f|_{f^{-1}(\Sigma-\Delta_{f})}:f^{-1}(\Sigma-\Delta_{f})\to\Sigma-\Delta_{f}$, let $a^{*}f|_{f^{-1}(\Sigma-\Delta_{f})}:a^{*}f^{-1}(\Sigma-\Delta_{f})\to[0,1]$ denote the pullback of $f|_{f^{-1}(\Sigma-\Delta_{f})}$ by a smooth loop $a:[0,1]\to\Sigma-\Delta_{f}$. Let $\Phi:[0,1]\times S_{g}\to a^{*}f^{-1}(\Sigma-\Delta_{f})$ be a trivialization map of $a^{*}f|_{f^{-1}(\Sigma-\Delta_{f})}$ with $\Phi(0,p)=(0,\Xi^{-1}(p))$ for all $p\in S_{g}$, and for each $t\in[0,1]$, define a map $\Phi_{t}:S_{g}\to f^{-1}(y)$ as $\Phi_{t}(p):=proj_{2}\circ\Phi(t,p)\thickspace(p\in S_{g})$, where $proj_{2}:[0,1]\times f^{-1}(\Sigma-\Delta_{f})\to f^{-1}(\Sigma-\Delta_{f})$ is the second projection. Then we get a well-defined homomorphism $\rho_{f}:\pi_{1}(\Sigma-\Delta_{f},y)\to\mathcal{M}(S_{g})$ which is defined by $\rho_{f}([a]):=[\Xi\circ\Phi_{1}]$ for all $a]\in\pi_{1}(\Sigma-\Delta_{f},y)$. This is called a monodromy representation of $f$.
@@
For $i=1,\cdots n$, we say that the homotopy class of $a_{i}$ is the meridian loop of $f$ with respect to $\mu_{i}$. It is well-known that $\rho_{f}([a_{i}])$ is equal to the Dehn twist about the vanishing cycle of $f$ determined by $\mu_{i}$. Here, we note a few things about this. First, consider the case of $f$ being orientable. In this case, the Dehn twist about the vanishing cycle determined by $\mu_{i}$ is right handed (resp. left handed) if the critical point of $f$ belonging to $\mathrm{Crit}(f)\cap f^{-1}(y_{i})$ is positive (resp. negative). Next, consider the case of $f$ being non-orientable. From Remark 3.2, notice that the Dehn twist about each vanishing cycle of $f$ can be defined. In this case, whether the Dehn twist about the vanishing cycle determined by $\mu_{i}$ is right handed or left handed depends on an orientation of a tubular neighborhood of the vanishing cycle because of the relation (4.1).
@@
It is easily seen that the domains of $\rho_{f\circ\pi}$ and $\eta\circ\rho_{f}$ are the same since $\Sigma-\Delta_{f\circ\pi}=\Sigma-\Delta_{f}$ holds. By applying Lemma 2.5, for a fixed diffeomorphism $\Xi:f^{-1}(y)\to N_{g}$, we get a unique orientation preserving diffeomorphism $\widetilde{\Xi}:(f\circ\pi)^{-1}(y)\to\Sigma_{g-1}$ such that the diagram (2.1) commutes. Take the pull-back $a^{*}(f\circ\pi)|_{(f\circ\pi)^{-1}(\Sigma-\Delta_{f})}:a^{*}(f\circ\pi)^{-1}(\Sigma-\Delta_{f})\to[0,1]$ of the fiber bundle $(f\circ\pi)|_{(f\circ\pi)^{-1}(\Sigma-\Delta_{f})}:(f\circ\pi)^{-1}(\Sigma-\Delta_{f})\to\Sigma-\Delta_{f}$ by $a$. Note that $(id_{[0,1]}\times\pi)|_{a^{*}(f\circ\pi)^{-1}(\Sigma-\Delta_{f})}:a^{*}(f\circ\pi)^{-1}(\Sigma-\Delta_{f})\to a^{*}f^{-1}(\Sigma-\Delta_{f})$ is an orientation double covering map since $a^{*}f^{-1}(\Sigma-\Delta_{f})$ is a submanifold of $0,1]\times f^{-1}(\Sigma-\Delta_{f})$ and $a^{*}(f\circ\pi)^{-1}(\Sigma-\Delta_{f})=\pi^{-1}(a^{*}f^{-1}(\Sigma-\Delta_{f}))$. Lemma 2.5 hence gives a unique orientation preserving diffeomorphism $\widetilde{\Phi}:[0,1]\times\Sigma_{g-1}\to a^{*}(f\circ\pi)^{-1}(\Sigma-\Delta_{f})$ for the diffeomorphism $\Phi:[0,1]\times N_{g}\to a^{*}f^{-1}(\Sigma-\Delta_{f})$ so that the upper rectangle of the diagram